1. suppose you have the following information for a given economy (the values are in billion Birr)


GDP  = 5000

Factor income from abroad  = 40

Factor payment to abroad = 15

Capital accumulation allowance (depreciation) = 450

Personal tax and non tax payments = 600

Social security contributions = 550

Government transfers to individuals = 755

Corporate profit = 300

Dividends = 120

Indirect business taxes = 400

Calculate :

  1. GNP 
  2. National income 
  3. disposable income
1
Expert's answer
2022-03-16T11:28:22-0400

Q311052

Factor income from abroad = 40

Factor payment to abroad = 15

Capital accumulation allowance (depreciation) = 450

Personal tax and non tax payments = 600

Social security contributions = 550

Government transfers to individuals = 755

Corporate profit = 300

Dividends = 120

Indirect business taxes = 400

Calculate :

  1. GNP 

Solution

GNP =GDP -Net income payable to the rest of the world

Net income payments to the rest of the world = factor payments to abroad - factor income for abroad.

= (15 - 40) = - 25

GNP = 5000 - (-25)

= 5025


2. National income 

Solution

National income = GNP - Depreciation

= 5025 - 450

= 4575

3. Disposable income

Solution

Disposable income = personal income - personal tax and non tax payments

personal income = National income - (social security contributions + government transfer to individuals - corporate profit + dividends - indirect taxes)

= 4575 - (550 + 255 - 300 +120 - 400)

= 4200

Therefore disposable income = 4200 - 600

= 3600